OpenStudy (anonymous):

find the equation of the line that is parelle to th eline 3y-x=9 and passes through (-2,1)?

OpenStudy (anonymous):

3y - x = 9 3y = x + 9 y = x / 3 + 3 Gradient = 1 / 3 Since the lines are parallel, the gradient is the same. Using (-2, 1): y = 1 + (x + 2) / 3 3y = 3 + x + 2 3y = x + 5

OpenStudy (anonymous):

3y = x + 5